# ChestX-Reasoner: Advancing Radiology Foundation Models with Reasoning through Step-by-Step Verification

### v1

@@ Abstract
Recent advances in reasoning-enhanced large language models (LLMs) and multimodal LLMs (MLLMs) have significantly improved performance in complex tasks, yet medical AI models often overlook the structured reasoning processes inherent in clinical practice. In this work, we present ChestX-Reasoner, a radiology diagnosis MLLM designed to leverage process supervision mined directly from clinical reports, reflecting the step-by-step reasoning followed by radiologists. We construct a large dataset by extracting and refining reasoning chains from routine radiology reports. Our two-stage training framework combines supervised fine-tuning and reinforcement learning guided by process rewards to better align model reasoning with clinical standards. We introduce RadRBench-CXR, a comprehensive benchmark featuring 59K visual question answering samples with 301K clinically validated reasoning steps, and propose RadRScore, a metric evaluating reasoning factuality, completeness, and effectiveness. ChestX-Reasoner outperforms existing medical and general-domain MLLMs in both diagnostic accuracy and reasoning ability, achieving 16%, 5.9%, and 18% improvements in reasoning ability compared to the best medical MLLM, the best general MLLM, and its base model, respectively, as well as 3.3%, 24%, and 27% improvements in outcome accuracy. All resources are open-sourced to facilitate further research in medical reasoning MLLMs.
@@ 1 Introduction
For dataset construction, we highlight that daily radiology reports are an invaluable resource for constructing high-quality reasoning chains, as they typically include doctors’ findings, and impressions. To this end, as illustrated in Fig. 1(b) and Fig. 2(a), we propose to construct diverse question-answer pairs based on impressions, and extracts clinically relevant observations in findings as the reasoning processes. We leverage tools like GPT-4o [15] to automate the labor-intensive tasks, for example, generating question-answer pairs, extracting clinical observations, and refining them into complete, clinically coherent reasoning chains. Compared to the collecting reasoning processes from large visual-language models or relying on manual annotation, extracting reasoning chains directly from textual reports is both scalable and highly reliable.
@@
For model evaluation, we curate a reliable reasoning benchmark on Chest X-rays, termed as RadRBench-CXR, that features detailed diagnostic reasoning processes. It includes 59K visual question answering (VQA) samples from three public datasets, namely, MIMIC-CXR [17], ChexPert [18], and MS-CXR-T [19]. The questions span diverse task types that require complex reasoning chains. Each sample is augmented with reasoning processes mined from its corresponding clinical report, comprising 301k clinically validated reasoning steps. Additionally, we present an evaluation metric, RadRScore, that assesses a model’s reasoning capabilities across three key dimensions: factuality (the correctness of generated reasoning), completeness (the thoroughness in covering clinical findings), and effectiveness (the necessity and relevance of diagnostic processes).
As a result, our proposed ChestX-Reasoner is a multimodal large language model with medical reasoning for Chest X-ray diagnosis. We compare it against three medical foundation models and five general-domain foundation models in terms of outcome accuracy and reasoning ability, using RadRScore on both RadRBench-CXR and the public CheXbench [20]. As demonstrated in Fig. 1(d), ChestX-Reasoner achieves significant improvements compared to all baselines. Specifically, it achieves 16%, 5.9%, and 18% improvements in reasoning ability compared to the best medical baseline, the best general baseline, and its base model (Qwen2VL-7B [21]), respectively, as well as 3.3%, 24%, and 27% improvements in outcome accuracy. All code, datasets, and models are open-sourced at ChestX-Reasoner, and we hope this work inspires further research into developing medical reasoning MLLMs.
@@ 2 Benchmarks
Figure 2: Reasoning Mining Procedures and RadRScore Calculation a. Procedures of mining reasoning. We first construct reasoning plans by prompting GPT-4o based on given QA pairs and clinical reports, then extract relative clinical observations from the report as answers for each plan to derive the reasoning steps, and finally, refine the logic to form coherent and complete reasoning chains. b. RadRScore calculation targeting on factuality (the correctness of generated reasoning), completeness (the thoroughness in covering clinical findings), and effectiveness (the necessity and relevance of diagnostic processes).
@@
We use $\text{obs}_{\text{model}}$, $\text{obs}_{\text{gt}}$, and $\text{obs}_{\text{report}}$ to respectively denote the extracted entity set from the model’s output, ground-truth reasoning, and corresponding report. Then, we calculate the values of following three dimensions, namely factuality, completeness, and effectiveness based on $\text{obs}_{\text{model}}$, $\text{obs}_{\text{gt}}$, and $\text{obs}_{\text{report}}$:
@@
To create a balanced and unbiased dataset, we incorporate a diverse range of diseases, organs, and findings while ensuring variety across samples. For the balancing strategy, we analyze the disease distribution in the original data and adjust the quantities, such that the most frequent disease does not exceed twice the least frequent ones, promoting fairness in the dataset.
Specifically, the benchmark comprises 1.2M answer-only VQA samples for training and 7.3K samples for testing. For a subset of these, we generate 59K reasoning-augmented VQA samples for training and 1K reasoning-augmented samples for testing. We apply strict quality control to the VQA samples with reasoning chains. Samples with a factuality score $R_{f}$ lower than 1 are filtered out, ensuring that only high-quality, factual reasoning chains are retained in our benchmark. Notably, prior to factuality filtering, the generated reasoning processes achieve a significantly higher factuality score ($R_{f}$) of 0.82, compared to GPT-4o (0.60) and Qwen2VL-72B (0.59), demonstrating the effectiveness of our method to produce high quality reasoning samples. The detailed cases of instruction design, and mined reasoning processes are shown in Supplementary 6.3.
@@ 3 Results: Reasoning Enhanced Medical VLM (ChestX-Reasoner)
The comparison includes three medical MLLMs—RadFM-14B [24], CheXagent-3B [20], and MedDr-40B [25], and five general MLLMs: DeepSeek-VL-7B [26], LLaVA-NeXT-8B [27], Qwen2VL-7B, Qwen2VL-72B [21], and GPT-4o [15]. The evaluation encompasses the five tasks explained in Sec. 2, namely, binary disease diagnosis, single disease diagnosis, multiple disease diagnosis, anomaly detection, and temporal comparison analysis.
Due to the lack of clinical reports in SIIM [23] and RSNA [22], no mined reasoning processes or process supervision were applied to samples from these datasets during our training pipeline. We use these two datasets for cross-center validation. Additionally, we provide an ablation study on the impact of different training strategies, underscoring the critical role of process supervision in achieving superior outcomes.
@@ 3.1 Comparison on Reasoning Ability
Factuality. As shown in the first column of Fig. 3(b), ChestX-Reasoner achieved the following scores: 0.660  (95%CI = 0.603–0.712) on binary disease diagnosis, 0.741  (95%CI = 0.672–0.802) on single disease diagnosis, 0.719  (95%CI = 0.691–0.746) on multiple disease diagnosis, 0.517  (95%CI = 0.466–0.566) on anomaly detection, and 0.715  (95%CI = 0.626–0.800) on temporal comparison analysis. These results outperform the best general MLLM, GPT-4o with average scores of 0.534, 0.617, 0.554, 0.460, and 0.532 on the five tasks, respectively. Additionally, ChestX-Reasoner also outperforms the best medical MLLM, MedDr-40B, which achieved average scores of 0.510, 0.443, 0.508, 0.348, and 0.589 on the tasks. These findings demonstrate that ChestX-Reasoner exhibits superior clinical factual correctness compared to other baselines.
Completeness. As shown in the second column of Fig. 3(b), ChestX-Reasoner achieved the following scores: 0.248  (95%CI = 0.213–0.281) on binary disease diagnosis, 0.438  (95%CI = 0.372–0.507) on single disease diagnosis, 0.503  (95%CI = 0.476–0.531) on multiple disease diagnosis, 0.298  (95%CI = 0.260–0.339) on anomaly detection, and 0.284  (95%CI = 0.221–0.347) on temporal comparison analysis. Among the baselines, general-domain MLLMs like GPT-4o and Qwen2VL-72B achieved comparable performance, except for the anomaly detection task. Qwen2VL-72B failed in anomaly detection due to the lack of given options. In contrast, medical-domain MLLMs performed significantly worse in this dimension. The best-performing medical MLLM, MedDr-40B, achieved completeness scores of only 0.155 on binary disease diagnosis, 0.171 on single disease diagnosis, 0.291 on multiple disease diagnosis, 0.156 on anomaly detection, and 0.082 on temporal comparison analysis. These results underscore the superior thoroughness of ChestX-Reasoner in covering clinical findings while highlighting the limitations of current medical-domain MLLMs.
Effectiveness. As shown in the third column of Fig. 3(b), ChestX-Reasoner achieved the following scores: 0.619  (95%CI = 0.563–0.674) on binary disease diagnosis, 0.690  (95%CI = 0.614–0.756) on single disease diagnosis, 0.547  (95%CI = 0.510–0.579) on multiple disease diagnosis, 0.399  (95%CI = 0.333–0.472) on anomaly detection, and 0.630  (95%CI = 0.537–0.728) on temporal comparison analysis. In this dimension, the medical-domain MLLM, MedDR-40B, achieves performance comparable to the best general-domain MLLM, GPT-4o, but still falls significantly short of ChestX-Reasoner. Our ChestX-Reasoner outperforms both GPT-4o and MedDR-40B, with improvements of 5.4% and 24.8%, respectively. These results underscore the superior necessity and relevance of the diagnostic processes generated by our model.
@@ 3.2 Comparison on Outcome Accuracy
Figure 5: Ablation studies on both reasoning ability and outcome accuracy of our trained models including ChestX-Reasoner-SFT(RO), ChestX-Reasoner-SFT(Both), ChestX-Reasoner-RL(O), ChestX-Reasoner-SFT(RO)+RL(O), ChestX-Reasoner (w.o. Progress Reward), and the final ChestX-Reasoner.
3.3 Training Strategy Ablation
@@
To address this gap, we conducted an ablation study (Fig. 5) comparing various strategies, resulting in models:
@@ 4 Discussion
LLMs can transform clinical reports into reasoning samples across various radiology tasks. The development of medical reasoning MLLMs is often constrained by the high cost and limited availability of high-quality, reliable multimodal reasoning datasets. In this work, we demonstrate that existing clinical reports, combined with effective prompt engineering, enable current LLMs to automatically generate diverse, well-structured reasoning-labeled training instances across a broad range of radiology tasks. Notably, our generated reasoning, based on clinical reports, achieves a significantly higher factuality score (the correctness of generated reasoning defined in RadRScore) of 0.82, substantially outperforming GPT-4o (0.60) and QwenVL-72B (0.59).
@@ 5.1 Preliminaries
Assuming dataset $\mathcal{D}=\{(\mathcal{X}_{i},\mathcal{Q}_{i},\mathcal{P}_{i},\mathcal{C}_{i},\mathcal{A}_{i})\}_{i=1}^{N}$, where $N$ denotes the number of samples, $\mathcal{X}_{i}$ represents the input images, $\mathcal{Q}_{i}$ denotes the associated text instruction, $\mathcal{P}_{i}$ refers to the mined reasoning process, $\mathcal{C}_{i}$ is the corresponding clinical report, and $\mathcal{A}_{i}$ is the ground-truth answer. Our objective is to develop a multimodal large language model (MLLM) that enhances Chest X-ray interpretation by improving both reasoning capability and answer accuracy on $\mathcal{D}$. Formally, the objective is to generate the desired output sentences given the input images $\mathcal{X}$ and the text instruction $\mathcal{Q}$:
@@ 5.2 ChestX-Reasoner
where $\epsilon$ and $\beta$ are hyperparameters controlling the clipping range and the strength of the KL divergence penalty, respectively. KL penalty constrains the updating calculated as: $D_{\text{KL}}\left(\pi_{{}_{\text{old}}}\parallel\pi_{\theta}\right)=\log\left(\frac{\pi_{{\text{old}}}(o_{i}|\mathcal{X},\mathcal{Q})}{\pi_{{\theta}}(o_{i}|\mathcal{X},\mathcal{Q})}\right).$ The advantage $A_{i}$ is computed using a group of rewards $\{r_{1},r_{2},\dots,r_{G}\}$ corresponding to the group outputs calculated as $A_{i}=\frac{r_{i}-\text{mean}(\{r_{1},r_{2},\dots,r_{G}\})}{\text{std}(\{r_{1},r_{2},\dots,r_{G}\})}.$ Samples from $\mathcal{D}_{A}$ use the outcome format reward and the outcome accuracy reward, whereas for samples from $\mathcal{D}_{R}$, an additional process reward is included, which are all defined as follows:
@@
where $\mathcal{C}$ is the clinical report, and $\mathcal{A}$ is ground truth answer. The distinction among these rewards lies in their supervision signals. From the equation, we can easily find that, outcome format reward and outcome accuracy reward are calculated based on only several tokens describing format or final answer compared to a specific format and ground truth answer $\mathcal{A}$. These two rewards are typically used in many prior work [2, 7, 6, 13, 14], as well as in ChestX-Reasoner-RL(O) and ChestX-Reasoner-SFT(RO)+RL(O), lacking supervision of reasoning processes. To incentivize high-quality reasoning, we preserves the robust reasoning supervision embedded in clinical reports via process reward, ensuring a more transparent and reliable reasoning process through the process factuality reward.
@@
All datasets used in this study are publicly accessible. This study utilized datasets that are all publicly accessible. The RadRBench-CXR evaluation benchmark, and the model weights of all stages and variants will be released before publication.
@@
The code used for experiments in this study will be made publicly available before publication. It includes the code to create reasoning processes from reports and the code to train the model ChestX-Reasoner. All the models will be hosted on HuggingFace (https://huggingface.co/) before publication.

### v2

@@ Abstract
Recent advances in reasoning-enhanced large language models (LLMs) and multimodal LLMs (MLLMs) have significantly improved performance in complex tasks, yet medical AI models often overlook the structured reasoning processes inherent in clinical practice. In this work, we present ChestX-Reasoner, a radiology diagnosis MLLM designed to leverage process supervision mined directly from clinical reports, reflecting the step-by-step reasoning followed by radiologists. We construct a large dataset by extracting and refining reasoning chains from routine radiology reports. Our two-stage training framework combines supervised fine-tuning and reinforcement learning guided by process rewards to better align model reasoning with clinical standards. We introduce RadRBench-CXR, a comprehensive benchmark featuring 59K visual question answering samples with 301K clinically validated reasoning steps, and propose RadRScore, a metric evaluating reasoning factuality, completeness, and effectiveness. ChestX-Reasoner outperforms existing medical and general-domain MLLMs in both diagnostic accuracy and reasoning ability, achieving 16%, 8.5%, and 18% improvements in reasoning ability compared to the best medical MLLM, the best general MLLM, and its base model, respectively, as well as 3.3%, 24%, and 27% improvements in outcome accuracy. All resources are open-sourced to facilitate further research in medical reasoning MLLMs.
@@ 1 Introduction
For dataset construction, we highlight that daily radiology reports are an invaluable resource for constructing high-quality reasoning chains, as they typically include doctors’ findings and impressions. To this end, as illustrated in Fig. 1(b) and Fig. 2(a), we propose to construct diverse question-answer pairs based on impressions, and extracts clinically relevant observations in findings as the reasoning processes. We leverage tools like GPT-4o [15] to automate the labor-intensive tasks, for example, generating question-answer pairs, extracting clinical observations, and refining them into complete, clinically coherent reasoning chains. Compared to collecting reasoning processes from large visual-language models or relying on manual annotation, extracting reasoning chains directly from textual reports is both scalable and highly reliable.
@@
For model evaluation, we curate a reliable reasoning benchmark on Chest X-rays, termed as RadRBench-CXR, that features detailed diagnostic reasoning processes. It includes 59K visual question answering (VQA) samples from three public datasets, namely, MIMIC-CXR [17], CheXpert [18], and MS-CXR-T [19]. The questions span diverse task types that require complex reasoning chains. Each sample is augmented with reasoning processes mined from its corresponding clinical report, comprising 301k clinically validated reasoning steps. As shown in Fig. 2(c), the mined reasoning processes demonstrate high quality upon review by radiologists, and we filtered out unsatisfactory samples to ensure a reliable evaluation. Additionally, we present an evaluation metric, RadRScore, that assesses a model’s reasoning capabilities across three key dimensions: factuality (the correctness of generated reasoning), completeness (the thoroughness in covering clinical findings), and effectiveness (the necessity and relevance of diagnostic processes).
As a result, our proposed ChestX-Reasoner is a multimodal large language model with medical reasoning for Chest X-ray diagnosis. We compare it against three medical foundation models and five general-domain foundation models in terms of outcome accuracy and reasoning ability, using RadRScore on both RadRBench-CXR and the public CheXbench [20]. As demonstrated in Fig. 1(d), ChestX-Reasoner achieves significant improvements compared to all baselines. Specifically, it achieves 16%, 8.5%, and 18% improvements in reasoning ability compared to the best medical baseline, the best general baseline, and its base model (Qwen2VL-7B [21]), respectively, as well as 3.3%, 24%, and 27% improvements in outcome accuracy. All code, datasets, and models are open-sourced at ChestX-Reasoner, and we hope this work inspires further research into developing medical reasoning MLLMs.
@@ 2 Benchmarks
Figure 2: Reasoning Mining Procedures and RadRScore Calculation a. Procedures of mining reasoning. We first construct reasoning plans by prompting GPT-4o based on given QA pairs and clinical reports, then extract relative clinical observations from the report as answers for each plan to derive the reasoning steps, and finally, refine the logic to form coherent and complete reasoning chains. b. RadRScore calculation targeting on factuality (the correctness of generated reasoning), completeness (the thoroughness in covering clinical findings), and effectiveness (the necessity and relevance of diagnostic processes). c. Human rating on the factuality, completeness, and effectiveness of our mined reasoning processes in the test set, where we use a five-point Likert scale ranging from -10 to 10, based on the provided question, ground truth answer, X-ray images, and clinical report.
@@
We use $\text{obs}_{\text{model}}$, $\text{obs}_{\text{gt}}$, and $\text{obs}_{\text{report}}$ to respectively denote the extracted entity set from the model’s output, ground-truth reasoning, and corresponding report. Then, we calculate the values from three dimensions, namely factuality, completeness, and effectiveness based on $\text{obs}_{\text{model}}$, $\text{obs}_{\text{gt}}$, and $\text{obs}_{\text{report}}$:
@@
To create a balanced and unbiased dataset, we analyze the disease distribution in the original data and adjust the quantities, such that the most frequent disease does not exceed twice the least frequent ones, promoting fairness in the dataset.
Specifically, the benchmark comprises 1.2M answer-only VQA samples for training and 7.3K samples for testing. For a subset of these, we generate 59K reasoning-augmented VQA samples for training and 1K reasoning-augmented samples for testing. We apply strict quality control to the VQA samples with reasoning chains. Samples with a factuality score $R_{f}$ lower than 1 are filtered out, ensuring that only high-quality, factual reasoning chains are retained in our benchmark. Notably, prior to factuality filtering, the generated reasoning processes achieve a significantly higher factuality score ($R_{f}$) of 0.82, compared to GPT-4o (0.60) and Qwen2VL-72B (0.59), demonstrating the effectiveness of our method to produce high quality reasoning samples. Additionally, to ensure the reliability of the reasoning evaluation, we conducted human verification, where the radiologists were provided with Chest X-ray images, corresponding reports, and the reasoning outputs generated by our method on the test samples. As shown in Fig. 2(c), they achieve average scores of factuality: 7.83, validity: 7.25, and completeness: 7.97 (on a scale from -10 to 10), showing sufficiently high quality. For the released test samples, any reasoning process with a score below 5 in any of the three dimensions was filtered out. The detailed cases of instruction design, and mined reasoning processes are shown in Supplementary 6.3.
@@ 3 Results: Reasoning Enhanced Medical VLM (ChestX-Reasoner)
The comparison includes three medical MLLMs—RadFM-14B [24], CheXagent-3B [20], and MedDr-40B [25], and five general MLLMs: DeepSeek-VL-7B [26], LLaVA-NeXT-8B [27], Qwen2VL-7B, Qwen2VL-72B [21], and GPT-4o [15]. The evaluation encompasses the five tasks explained in Sec. 2, namely, binary disease diagnosis, single disease diagnosis, multiple disease diagnosis, anomaly detection, and temporal comparison analysis. Due to the lack of clinical reports in SIIM [23] and RSNA [22], no mined reasoning processes or process supervision were applied to samples from these datasets during our training pipeline. We use these two datasets for cross-center validation. Additionally, we provide an ablation study on the impact of different training strategies, underscoring the critical role of process supervision in achieving superior outcomes. We also provide visualizations of four cases in Supplementary Figs. 1, 2, 3, and 4, including the VQA questions, extracted reasoning processes, and model outputs from our ChestX-Reasoner, GPT-4o, and Qwen2VL-72B. Please refer to these visualizations to better understand the effectiveness of our model.
@@ 3.1 Comparison on Reasoning Ability
Factuality. As shown in the first column of Fig. 3(b), ChestX-Reasoner achieved the following scores: 0.682  (95%CI = 0.646–0.710) on binary disease diagnosis, 0.751  (95%CI = 0.722–0.782) on single disease diagnosis, 0.778  (95%CI = 0.753–0.800) on multiple disease diagnosis, 0.588  (95%CI = 0.560–0.615) on anomaly detection, and 0.706  (95%CI = 0.617–0.790) on temporal comparison analysis. These results outperform the best general MLLM, GPT-4o with average scores of 0.534, 0.617, 0.554, 0.460, and 0.532 on the five tasks, respectively. Additionally, ChestX-Reasoner also outperforms the best medical MLLM, MedDr-40B, which achieved average scores of 0.570, 0.633, 0.567, 0.471, and 0.532 on these tasks. These findings demonstrate that ChestX-Reasoner exhibits superior clinical factual correctness compared to other baselines.
Completeness. As shown in the second column of Fig. 3(b), ChestX-Reasoner achieved the following scores: 0.240  (95%CI = 0.218–0.262) on binary disease diagnosis, 0.430  (95%CI = 0.403–0.458) on single disease diagnosis, 0.520  (95%CI = 0.497–0.541) on multiple disease diagnosis, 0.282  (95%CI = 0.262–0.301) on anomaly detection, and 0.259  (95%CI = 0.199–0.323) on temporal comparison analysis. Among the baselines, general-domain MLLMs like GPT-4o and Qwen2VL-72B achieved comparable performance, except for the anomaly detection task. Qwen2VL-72B failed in anomaly detection due to the lack of given options. In contrast, medical-domain MLLMs performed significantly worse in this dimension. The best-performing medical MLLM, MedDr-40B, achieved completeness scores of only 0.155 on binary disease diagnosis, 0.171 on single disease diagnosis, 0.291 on multiple disease diagnosis, 0.156 on anomaly detection, and 0.082 on temporal comparison analysis. These results underscore the superior thoroughness of ChestX-Reasoner in covering clinical findings while highlighting the limitations of current medical-domain MLLMs.
Effectiveness. As shown in the third column of Fig. 3(b), ChestX-Reasoner achieved the following scores: 0.661  (95%CI = 0.627–0.695) on binary disease diagnosis, 0.688  (95%CI = 0.649–0.727) on single disease diagnosis, 0.597  (95%CI = 0.565–0.627) on multiple disease diagnosis, 0.415  (95%CI = 0.380-0.450) on anomaly detection, and 0.698  (95%CI = 0.600–0.796) on temporal comparison analysis. In this dimension, the medical-domain MLLM, MedDr-40B, achieves performance comparable to the best general-domain MLLM, GPT-4o, but still falls significantly short of ChestX-Reasoner. Our ChestX-Reasoner outperforms both GPT-4o and MedDr-40B, with improvements of 5.8% and 34.0%, respectively. These results underscore the superior necessity and relevance of the diagnostic processes generated by our model.
@@ 3.2 Comparison on Outcome Accuracy
Figure 5: Ablation studies on both reasoning ability and outcome accuracy of our trained models including ChestX-Reasoner-SFT(RO), ChestX-Reasoner-SFT(Both), ChestX-Reasoner-RL(O), ChestX-Reasoner-SFT(RO)+RL(O), ChestX-Reasoner (w.o. Process Reward), and the final ChestX-Reasoner.
3.3 Ablation on Training Strategy
@@
To bridge this gap, we conducted an ablation study (Fig. 5) comparing various strategies, resulting in models:
@@ 4 Discussion
LLMs can transform clinical reports into reasoning samples across various radiology tasks. The development of medical reasoning MLLMs is often hindered by the scarcity and high cost of high-quality, reliable multimodal reasoning datasets. In this work, we show that existing clinical reports, combined with effective prompt engineering, can enable LLMs to automatically generate diverse, well-structured reasoning-labeled training instances across a wide range of radiology tasks. By mining reasoning processes from clinical reports, our generated reasoning achieves a significantly higher factuality score-defined by RadRScore, which measures the correctness of generated reasoning-reaching 0.82, compared to GPT-4o (0.60) and QwenVL-72B (0.59). Additionally, human evaluations conducted by radiologists (Fig. 2(c)) reveal strong satisfaction with the generated reasoning, with average scores of factuality: 7.83, validity: 7.25, and completeness: 7.97 (on a scale from -10 to 10). These results highlight the potential of leveraging LLMs to convert clinical reports into high-quality reasoning samples, offering promising implications for broader applications in medical AI and beyond.
@@ 5.1 Preliminaries
Assuming dataset $\mathcal{D}=\{(\mathcal{X}_{i},\mathcal{Q}_{i},\mathcal{P}_{i},\mathcal{C}_{i},\mathcal{A}_{i})\}_{i=1}^{N}$, where $N$ denotes the number of samples, $\mathcal{X}_{i}$ represents the input images, $\mathcal{Q}_{i}$ denotes the associated text instruction, $\mathcal{P}_{i}$ refers to the mined reasoning processes, $\mathcal{C}_{i}$ is the corresponding clinical report, and $\mathcal{A}_{i}$ is the ground-truth answer. Our objective is to develop a multimodal large language model (MLLM) that enhances Chest X-ray interpretation by improving both reasoning capability and outcome accuracy on $\mathcal{D}$. Formally, the objective is to generate the desired output sentences given the input images $\mathcal{X}$ and the text instruction $\mathcal{Q}$:
@@ 5.2 ChestX-Reasoner
where $\epsilon$ and $\beta$ are hyperparameters controlling the clipping range and the strength of the KL divergence penalty, respectively. KL penalty constrains the updating calculated as: $D_{\text{KL}}\left(\pi_{{}_{\text{old}}}\parallel\pi_{\theta}\right)=\log\left(\frac{\pi_{{\text{old}}}(o_{i}|\mathcal{X},\mathcal{Q})}{\pi_{{\theta}}(o_{i}|\mathcal{X},\mathcal{Q})}\right).$ The advantage $A_{i}$ is computed using a group of rewards $\{r_{1},r_{2},\dots,r_{G}\}$ corresponding to the group outputs calculated as $A_{i}=\frac{r_{i}-\text{mean}(\{r_{1},r_{2},\dots,r_{G}\})}{\text{std}(\{r_{1},r_{2},\dots,r_{G}\})}.$ Samples from $\mathcal{D}_{A}$ use the outcome format reward and the outcome accuracy reward, whereas for samples from $\mathcal{D}_{R}$, an additional process factuality reward is included, which are all defined as follows:
@@
where $\mathcal{C}$ is the clinical report, and $\mathcal{A}$ is ground truth answer. The distinction among these rewards lies in their supervision signals. From the equation, we can easily find that, outcome format reward and outcome accuracy reward are calculated based on only several tokens describing format or final answer compared to a specific format and ground truth answer $\mathcal{A}$. These two rewards are typically used in many prior work [2, 7, 6, 13, 14], as well as in ChestX-Reasoner-RL(O) and ChestX-Reasoner-SFT(RO)+RL(O), lacking supervision of reasoning processes. To incentivize high-quality reasoning, we preserves the robust reasoning supervision embedded in clinical reports via process reward, ensuring a more transparent and reliable reasoning process.
@@
All datasets used in this study are publicly accessible. This study utilized datasets that are all publicly accessible. The RadRBench-CXR evaluation benchmark, and the model weights of all stages and variants are released in ChestX-Reasoner-Code.
@@
The code used for experiments in this study is released in ChestX-Reasoner-Code. It includes the code to create reasoning processes from reports and the code to train the model ChestX-Reasoner. We release the model weights in ChestX-Reasoner-Model and other variants will be hosted on HuggingFace (https://huggingface.co/) before publication.